# Six decades of the FitzHugh-Nagumo model: A guide through its spatio-temporal dynamics and influence across disciplines

### v3

@@ I.6 Outline
In this comprehensive review, we aim to offer an in-depth exploration of the diverse dynamical behaviors encapsulated within the FitzHugh-Nagumo (FHN) model. The widespread adoption of the FHN model across physics and biology can be attributed to the model’s remarkable versatility in capturing a wide array of dynamical phenomena while maintaining a relatively simple mathematical formulation.
Our review is structured around delineating the most prominent dynamical behaviors observed within the FHN model. We categorize our analysis into three primary sections: i) examining the foundational FHN model, characterized by a system of two nonlinear coupled ordinary differential equations (ODEs) [Eq. (8)]; ii) studying the diffusively coupled FHN model, which introduces spatial coupling through diffusion [Eq. (10)]; and iii) exploring discretely coupled FHN equations [Eq. (11)]. In each section, we complement our discussion of observed dynamics with thorough stability analyses and bifurcation studies. This approach empowers readers to navigate the parameter space effectively, enabling them to target specific dynamical regimes of interest.
@@
Figure 3: Overview of FitzHugh-Nagumo model applications This table organizes selected literature related to different formulations of the FitzHugh-Nagumo equations, as outlined in Eq. (8), Eq. (10), and Eq. (11). Each listed study is categorized by its application domain and primary focus. It is important to note that, alongside the conventional FitzHugh-Nagumo models, certain studies incorporate variations like piecewise linear models or three-dimensional adaptations to meet particular investigative requirements.
@@ II Dynamics in the FitzHugh-Nagumo model
Figure 4: Fixed points in the FHN model. A. The nullclines on the phase space ($\varepsilon=0.01$) and the effect of the parameters $a$ and $b$ on them are shown. B. The saddle node bifurcation in parameter space ($\varepsilon=0.01$) is shown and the mono- (blue) and tri-valued (red) regions is differentiated. SN indicates the saddle node bifurcations and their intersection in the cusp (C) bifurcation.
The basic FHN model [Eq. (8)] has been studied extensively because, despite its apparent simplicity, it exhibits complex dynamics and bifurcations. Here, we will limit ourselves to providing an overview of the most widely used dynamical regimes, such as monostability, multistability, relaxation oscillations, and excitability. For an in-depth bifurcation analysis, the reader is referred to the comprehensive work by Rocsoreanu et al. Rocsoreanu et al. (2012).
@@ II.1 Stationary solutions
The simplest solutions of Eq. (8) are the stationary states, also known as fixed points (FP) or equilibrium points. These points satisfy the condition where the time derivatives of both variables are zero [i.e., $(\dot{u}_{e},\dot{v}_{e})=(0,0)$], leading to a set of equations that define the so-called nullclines in the $(u,v)$ phase plane (Fig. 4A):
$\begin{array}[]{ @{} l @{} }0=-u^{3}+u-v,\\[3.0pt] 0=\varepsilon(u-bv+a).\end{array}$ (12)
At these nullclines, one of the variables remains constant over time: either $u$ (orange nullcline) or $v$ (green nullcline) does not change in time. The intersections of the nullclines represent the stationary solutions $(u,v)=(u_{e},v_{e})$, which are solutions to a cubic equation in terms of one of the variables ($u_{e}$):
$bu_{e}^{3}+(1-b)u_{e}+a=0,$ (13)
with $v_{e}=(u_{e}+a)/b.$ This cubic equation typically does not have a straightforward analytical solution, but its roots can indicate the existence of either one or three coexisting stationary states, depending on the system parameters. In the latter case, we label those solutions $u_{e}^{b}$, $u_{e}^{m}$, and $u_{e}^{t}$. As a result, the parameter space is divided into regions of monostability and multistability, delineated by fold lines that correspond to the conditions where the system transitions between having a single fixed point and multiple fixed points. These transitions are associated with saddle-node (SN) bifurcations, also called fold bifurcations, and the specific conditions under which they occur depend on the system parameters. These folds are located at $(u,a)=(u_{f}^{\pm},a_{f}^{\pm})$ where
$u_{f}^{\pm}=\pm\sqrt{\frac{b-1}{3b}},$ $a_{f}^{\pm}=\frac{2}{3}(1-b)\sqrt{\dfrac{b-1}{3b}},$ (14)
which are defined for $b\in(-\infty,0)\cup[1,\infty)$. The fixed points are triple-valued when $u_{f}^{-}<u_{e}<u_{f}^{+}$, and single-valued otherwise. The presence of these saddle-node bifurcations and their implications for the system’s dynamics are illustrated in Fig. 4A, which shows how variations in the parameters affect the number and positions of the stationary states.
Figure  4B further explores the parameter space ($a,b$), highlighting the regions associated with different dynamical behaviors and the critical points where these behaviors change. The fold lines $a=a_{f}^{\pm}(b)$ define the regions of monostability (blue - 1 FP) and multistability (red - 3 FP). For $b<0$, the system always shows a region of multistability. For $b>0$, multistability arises above the cusp bifurcation occurring at $b=1$ (C) where the fold points are created.
@@ II.2 Linear stability of the stationary solutions
To determine the linear stability of the fixed points, we introduce small perturbations of the form $(u,v)=(u_{e},v_{e})+\epsilon(\xi_{u},\xi_{v})e^{\sigma t}+c.c.$, with $\epsilon\ll 1$ Guckenheimer and Holmes (1983). A fixed point is considered stable if ${\rm Re}[\sigma]<0$, and unstable if ${\rm Re}[\sigma]>0$. The condition ${\rm Re}[\sigma]=0$ thus marks the occurence of a local bifurcation. To obtain $\sigma$, we compute the Jacobian matrix $J$ associated with the linearization of Eq. (8) around the equilibrium points and solve the linear eigenvalue problem
$(J-\sigma I_{2\times 2})\begin{pmatrix}\xi_{u}\\ \xi_{v}\end{pmatrix}=\begin{pmatrix}0\\ 0\end{pmatrix},$ $J\equiv\begin{pmatrix}1-3u_{e}^{2}&-1\\ \varepsilon&-\varepsilon b\end{pmatrix}.$ (15)
@@
$\begin{array}[]{ll}{\rm Det}(J)=\varepsilon(3bu_{e}^{2}-b+1),\\[3.0pt] {\rm Tr}(J)=1-3u_{e}^{2}-\varepsilon b.\end{array}$ (17)
Depending on the eigenvalues, two types of bifurcations may occur. If ${\rm Re}[\sigma]=0$ and ${\rm Im}[\sigma]=0$, the system undergoes a saddle-node bifurcation [Eq. (14)]. If ${\rm Re}[\sigma]=0$ and ${\rm Im}[\sigma]\neq 0$, a Hopf bifurcation occurs, leading to self-sustained oscillations. These conditions for limit cycle oscilations imply that ${\rm Tr}(J)=0$ and ${\rm Det}(J)>0$ which gives the following Hopf bifurcation point:
@@
$\sigma=\pm i\sqrt{{\rm Det}(J)}=\pm i\sqrt{\varepsilon(1-\varepsilon b^{2})},$ (19)
which provides the Hopf existence condition $b>1/\sqrt{\varepsilon}$.
@@ II.3 Dynamical regimes
Relaxation oscillations. Characterized by limit cycles, i.e., closed-orbit attractors in the phase space, which correspond to oscillations operating on distinct time scales—one fast and the other slow. An example of this nonlinear oscillation can be observed in Fig. 5A.1 (top) for parameters $a=0$ and $b=0.5$, marked by $\bullet$ in Fig. 5B. In the phase diagram (Fig. 5A, bottom), the oscillation traces a loop around an unstable equilibrium, the point where nullclines intersect. This oscillatory behavior is crucial for modeling biological rhythms, including mitotic Gelens et al. (2014), calcium Sneyd et al. (2017a) and cardiac oscillators Franzone et al. (2006), underscoring the FHN model’s applicability in simulating fast transitions between active and quiescent states in various biological systems.
@@
Excitability. This behavior is intrinsically linked to the oscillatory phenomena mentioned above and pertains to the system’s response to perturbation to a fixed point. The magnitude of the perturbation determines the system’s response. In the FHN framework, excitability manifests at parameter settings of $a=0.1$ and $b=1.5$, as illustrated in Fig. 5A.2. Minor perturbations result in a swift return to the baseline state, depicted in Fig. 5B with a light blue hue. Conversely, large perturbations drive the system along a nontrivial trajectory (excursion) in the phase space before resettling to the resting state, highlighted by a bold blue line in Fig. 5A.2. This trajectory, associated with a ’spike’ in temporal dynamics, is invariant to the perturbation’s specifics. Notably, excitability is the most used regime in neuroscience and it can reproduce neuronal spikes in the non-coupled cases Parra-Rivas et al. (2023a); FitzHugh (1961); Izhikevich (2007a) or in coupled cases Nagumo et al. (1962).
Bistability. A nonlinear system exhibits bistability when two stable fixed points coexist for the same set of parameters. This less frequently examined behavior within our model is depicted in Fig. 5A.3 for parameters $a=0$ and $b=2$. The phase space diagram illustrates how different initial conditions converge to separate stable fixed points, located where the nullclines intersect. Bistability contributes to the emergence of phenomena such as relaxation oscillations, excitability, and hysteresis Boissonade and De Kepper (1980); Rosanov (2002); Hock et al. (1993), vital for the model’s relevance and applicability.
@@
Figure 5: Dynamics and bifurcations of the FHN model for $\varepsilon=0.01$. A. Representative time series and phase space dynamics showing relaxation oscillations ($b=0.5$), excitability ($b=1.5$ and $a=0.1$), and bistability ($b=2$). B. Analytically derived Hopf and saddle node bifurcations, along with homoclinic bifurcations determined numerically, delineating various dynamical regions in the ($a,b$) parameter space. An expanded view of a specific area, marked by a black rectangle, is shown on the right. Black dots correspond to parameters used in A, and dashed lines mark sections analyzed in C. C Bifurcation diagrams of three representative regions. The diagram on the left (C.1) hints at an imminent homoclinic bifurcation due to the close proximity of the limit cycle to the saddle-node. The right diagram (C.3), on the other hand, reveals a shift from stability to instability, lacking an associated limit cycle, suggesting a global bifurcation has taken place.
@@
The complexity of the dynamics within the FHN model necessitates a more detailed analysis, as depicted in Fig. 5B. By examining cross-sections of this figure at constant $b$ values, we can generate 1D bifurcation diagrams, such as those shown in Fig. 5C, plotting the variable $u$ against $a$ for selected $b$ values.
@@
At $b=0.5$ (Fig. 5C.2), a single fixed point that varies monotonically with $a$ is present. This point undergoes Hopf bifurcations at H+ and H-, as indicated in Fig. 5B, leading to a distribution of stability shown in Fig. 5C.2, where solid and dashed lines denote stable and unstable equilibria, respectively. Unstable equilibria evolve into relaxation oscillations as illustrated in Fig. 5A.1, with limit cycles forming and dissapearing at H+ and H-. Through path continuation techniques, we can trace these cycles and assess their stability Doedel et al. (1991). In this parameter regime, oscillations are stable and exhibit amplitude variations as seen in Fig. 5C.2 (blue solid line). Close to H+ and H- the oscillations undergo a canard explosion characterized by the abrupt increase of the cycle amplitude for small changes of $a$ Bold et al. (2003).
@@
As $b$ increases, a cusp bifurcation at $(a,b)=(0,2)$ gives rise to three new equilibria: $u_{e}^{b}$, $u_{e}^{m}$, and $u_{e}^{t}$. Fig. 5C.3 for $b=2$ illustrates this, where H+ and H-, previously associated with opposite $a$ values, now occur close to SNr and SNl, respectively. This diagram corresponds to the most right vertical dashed line shown in the close-up view of Fig. 5B. Here the top and bottom equilibrium branches ($u_{e}^{t,b}$) are stable nodes and therefore yield a bistable regime. The former stable limit cycle has vanished due to a Bautin bifurcation occurring below the crossing point between H+ and H- at $b\approx 1.5$ (not shown here). At smaller $\varepsilon$ values, the Bautin point occurs near H+ and H-, thoroughly explored in Rocsoreanu’s work Rocsoreanu et al. (2012). Additionally, an unstable limit cycle, emerging from the inverted H+ and H-, quickly vanishes due to another homoclinic bifurcation occurring close to the Hopf bifurcation. For a deeper analysis of these phenomena at larger $\varepsilon$ values, Rocsoreanu et al.’s book offers comprehensive insights Rocsoreanu et al. (2012). Adjacent to this, type II excitability emerges, akin to that shown in Fig. 5A.2, characterized by its proximity to Hopf/Bautin bifurcations, facilitated by the time scale separation and the influence of the former limit cycle in the phase space. As mentioned in the introduction, this type of excitability is characterized by the absence of a clearly defined threshold.
@@
For $b<0$, the system’s bifurcation scenario changes significantly. An illustration for $b=-0.5$ is provided in Fig. 5C.1, where three fixed points persist, albeit with an opposite slope compared to the scenario in Fig. 5C.3. Unlike the bistable regime observed for $b>1$, the top and bottom fixed points ($u_{e}^{t,b}$) become unstable saddle nodes. Additionally, Hopf bifurcations that initially occurred at $u_{e}^{t,b}$ now emerge in the middle branch $u_{e}^{m}$, rendering it stable between SNl and H-, as well as between H+ and SNr. The relaxation oscillations stemming from H± also experience a canard explosion (see blue curve). Further decreasing $b$ leads to a homoclinic bifurcation (Hom.), indicated by the green line in Fig. 5B, where the oscillation period diverges, potentially exhibiting type-I excitability Izhikevich (2007b); Yelo-Sarrión et al. (2022). This aspect remains largely unexplored within the FHN model.
@@
    Monostable regime with a single equilibrium. Defined by a single stable node for $b>0$, this area splits into sectors I.1 and I.2 based on $a$’s positive and negative values, respectively. Both sectors exhibit type I excitability near the Hopf bifurcation, particularly in the presence of canard explosions Izhikevich (2007b).
@@
    Bistable regime. Situated between the saddle-node bifurcations SNl,r, this zone’s boundaries are demarcated by H+ for $a>0$ and H- for $a<0$, featuring three equilibria: two stable nodes ($u_{e}^{b,t}$) and one saddle point ($u_{e}^{m}$).
@@
    Oscillatory regime. This area is known for relaxation oscillations and is further divided into three subregions. In III.1 (for $b<0$), it is confined by Hopf bifurcations H± and the homoclinic bifurcation Hom.; in III.2 (for $b>0$), oscillations are bounded by H+,- and cease after crossing the Bautin bifurcation, which occurs near the Hopf bifurcation before the intersection of H+ and H-. Notably, canard explosions are observable near the Hopf bifurcations in each subregion.
@@
    Monostable regime within a tri-valued region. For $b<0$, this category includes two monostable regimes where a single stable node ($u_{e}^{m}$) coexists with two saddle points ($u_{e}^{b,t}$). These settings are located between SNl and H+ (IV.1) and between H- and SNr (IV.2). For $b>0$, a shift from bistable to monostable occurs due to H+ (for $a>0$) and H- (for $a<0$), leading to thin monostable strips flanked by H+ and SNr (for $a<0$) or by H- and SNl (for $a>0$).
@@
    Divergent dynamics regime. Characterized by divergent dynamics due to the lack of an attractor, this regime is prominent for $b<0$ and includes three main divergent sectors: V.1, below SNl, with only the saddle point $u_{e}^{b}$; V.2, above SNr, with just the saddle point $u_{e}^{t}$; and V.3, within the bounds of Hom., H+, and H-, where the unstable node $u_{e}^{m}$ coexists with saddles $u_{e}^{b,t}$. Proximity to the homoclinic bifurcation may induce type I excitability Izhikevich (2007b).
Figure 6: Time scale separation influences the shape and amplitude of the limit cycle. A. Effect of $\varepsilon$ on the dynamical behavior in phase space and time series for parameters $(a,b)=(0,0.5)$. Increasing $\varepsilon$ leads to less time scale separation and more sinusoidal oscillations. B. Relaxation oscillations disappear for larger values of $a$ and $\varepsilon$. The oscillation amplitude $A(a,\varepsilon)$ becomes dependent on the $a$ parameter, evidenced by the density plot in the left panel and the dispersion of scatter points in the right panel. Furthermore, an inverse relationship between amplitude and $\varepsilon$ is noted, with a decreasing trend in amplitude as $\varepsilon$ increases, indicated by the median and the 50% interquartile range values.
Previously introduced yet not fully explained phenomena, including canard explosions, homoclinic bifurcations (Hom.), and Bautin bifrucations (fold of cycles), warrant further exploration to better understand their roles in the dynamics observed.
A canard explosion refers to a quick increase in the limit cycle’s amplitude as the control parameter increases, generating a cycle that occupies a substantial portion of the phase space, instead of a small amplitude oscillation Wechselberger (2007); Benoit (1990). This phenomenon, illustrated in Figures 5A.1, 5C.1, and 5C.2, is also a characteristic of the Hodgkin-Huxley model Hasan et al. (2018); Rubin and Wechselberger (2007); Moehlis (2006); Ginoux and Llibre (2015) and results from the separation of time scales, a property preserved in the FHN model Rotstein et al. (2012); Parra-Rivas et al. (2023a); Ginoux and Llibre (2015); Rocsoreanu et al. (2012). The resilience of canard cycle amplitudes to parameter perturbations underscores their significance, with the presence of canards signaling type II excitability — a persistence of trajectory behavior after a limit cycle disappears.

### v4

@@ I.6 Outline
In this review, we aim to offer an in-depth exploration of the diverse dynamical behaviors encapsulated within the FHN model. The widespread adoption of the FHN model across physics and biology can be attributed to the model’s remarkable versatility in capturing a wide array of dynamical phenomena while maintaining a relatively simple mathematical formulation.
Our review is structured around delineating the most prominent dynamical behaviors observed within the FHN model. We categorize our analysis into three primary sections: (i) examining the foundational FHN model, characterized by a system of two nonlinear coupled ordinary differential equations (ODEs) [Eq. 8]; (ii) studying the diffusively coupled FHN model, which introduces spatial coupling through diffusion [Eq. 10]; and (iii) exploring discretely coupled FHN equations [Eq. 11]. In each section, we complement our discussion of observed dynamics with thorough stability analyses and bifurcation studies. This approach allows readers to navigate the parameter space effectively, enabling them to target specific dynamical regimes of interest.
@@ II Dynamics in the FitzHugh-Nagumo model
Figure 4: Fixed points in the FHN model. A. The nullclines in the ($u,v$) phase space ($\varepsilon=0.01$) and the effect of the parameters $a$ and $b$ on them are shown. B. The saddle node bifurcation in parameter space ($\varepsilon=0.01$) is shown and the mono- (1 FP) and tri-valued (3 FP) regions are differentiated. SN indicates the saddle node bifurcations and their intersection in the cusp (C) bifurcation.
The basic FHN model [Eq. 8] has been studied extensively because, despite its apparent simplicity, it exhibits complex dynamics and bifurcations. Here, we will limit ourselves to providing an overview of the most widely used dynamical regimes, such as monostability, multistability, relaxation oscillations, and excitability. For an in-depth bifurcation analysis, the reader is referred to the comprehensive work by Rocsoreanu et al. Rocsoreanu et al. (2012).
@@ II.1 Stationary solutions
The simplest solutions of Eq. 8 are the stationary states, also known as fixed points (FP) or equilibrium points. These points satisfy the condition where the time derivatives of both variables are zero [i.e., $(u_{\text{FP}_{t}},v_{\text{FP}_{t}})=(0,0)$], leading to a set of equations that define the so-called nullclines in the $(u,v)$ phase plane (Fig. 4A):
$\begin{array}[]{ @{} l @{} }0=-u^{3}+u-v,\vspace{3pt}\\ 0=\varepsilon(u-bv+a).\end{array}$ (12)
At these nullclines, one of the variables remains constant over time: either $u$ or $v$ does not change in time. The intersections of the nullclines represent the stationary solutions $(u,v)=(u_{\text{FP}},v_{\text{FP}})$, which are solutions to a cubic equation in terms of one of the variables ($u_{\text{FP}}$):
$bu_{\text{FP}}^{3}+(1-b)u_{\text{FP}}+a=0,$ (13)
with $v_{\text{FP}}=(u_{\text{FP}}+a)/b$. This cubic equation typically does not have a straightforward analytical solution, but its roots can indicate the existence of either one or three coexisting stationary states, depending on the system parameters. In the latter case, we label those solutions $u_{\text{FP}}^{b}$, $u_{\text{FP}}^{m}$, and $u_{\text{FP}}^{t}$. As a result, the parameter space is divided into regions of monostability and multistability, delineated by fold lines that correspond to the conditions where the system transitions between having a single fixed point and multiple fixed points. These transitions are associated with saddle-node (SN) bifurcations, also called fold bifurcations, and the specific conditions under which they occur depend on the system parameters. These folds are located at $(u,a)=(u_{\text{SN}}^{\pm},a_{\text{SN}}^{\pm})$ where
$u_{\text{SN}}^{\pm}=\pm\sqrt{\frac{b-1}{3b}},a_{\text{SN}}^{\pm}=\frac{2}{3}(1-b)\sqrt{\dfrac{b-1}{3b}},$ (14)
which are defined for $b\in(-\infty,0)\cup[1,\infty)$. The fixed points are triple-valued when $u_{\text{SN}}^{-}<u_{\text{FP}}<u_{\text{SN}}^{+}$, and single-valued otherwise. The presence of these saddle-node bifurcations and their implications for the system’s dynamics are illustrated in Fig. 4A, which shows how variations in the parameters affect the number and positions of the stationary states.
Figure 4B further explores the parameter space ($a,b$), highlighting the regions associated with different dynamical behaviors and the critical points where these behaviors change. The fold lines $a=a_{\text{SN}}^{\pm}(b)$ define the regions of monostability (1 FP) and multistability (3 FP). For $b<0$, the system always shows a region of multistability. For $b>0$, multistability arises above the cusp bifurcation occurring at $b=1$ (C) where the fold points are created.
@@ II.2 Linear stability of the stationary solutions
To determine the linear stability of the fixed points, we introduce small perturbations of the form $(u,v)=(u_{\text{FP}},v_{\text{FP}})+\epsilon(\xi_{u},\xi_{v})e^{\sigma t}+c.c$., with $\epsilon\ll 1$ Guckenheimer and Holmes (1983). A fixed point is considered stable if ${\rm Re}[\sigma]<0$, and unstable if ${\rm Re}[\sigma]>0$. The condition ${\rm Re}[\sigma]=0$ thus marks the occurrence of a local bifurcation. To obtain $\sigma$, we compute the Jacobian matrix $J$ associated with the linearization of Eq. 8 around the equilibrium points and solve the linear eigenvalue problem
$(J-\sigma I_{2\times 2})\begin{pmatrix}\xi_{u}\\ \xi_{v}\end{pmatrix}=\begin{pmatrix}0\\ 0\end{pmatrix},J\equiv\begin{pmatrix}1-3u_{\text{FP}}^{2}&-1\\ \varepsilon&-\varepsilon b\end{pmatrix}.$ (15)
@@
$\begin{array}[]{ll}{\rm Det}(J)=\varepsilon(3bu_{\text{FP}}^{2}-b+1),\vspace{3pt}\\ {\rm Tr}(J)=1-3u_{\text{FP}}^{2}-\varepsilon b.\end{array}$ (17)
Depending on the eigenvalues, two types of bifurcations may occur. If ${\rm Re}[\sigma]=0$ and ${\rm Im}[\sigma]=0$, the system undergoes a saddle-node bifurcation [Eq. 14]. If ${\rm Re}[\sigma]=0$ and ${\rm Im}[\sigma]\neq 0$, a Hopf bifurcation occurs, leading to self-sustained oscillations. These conditions for limit cycle oscillations imply that ${\rm Tr}(J)=0$ and ${\rm Det}(J)>0$ which gives the following Hopf bifurcation point:
@@
$\sigma=\pm i\sqrt{{\rm Det}(J)}=\pm i\sqrt{\varepsilon(1-\varepsilon b^{2})}.$ (19)
Equations (18) and (19) provide the conditions for the Hopf bifurcation. The zero trace condition implies $b\leq 1/\varepsilon$, whereas the positive determinant condition requires $b<1/\sqrt{\varepsilon}$.
@@ II.3 Dynamical regimes
Relaxation oscillations. Characterized by limit cycles, i.e., closed-orbit attractors in the phase space, which correspond to oscillations operating on distinct time scales – one fast and the other slow. An example of this nonlinear oscillation can be observed in Fig. 5A.1 (top) for parameters $a=0$ and $b=0.5$, marked by $\bullet$ in Fig. 5B. In the phase diagram (Fig. 5A, bottom), the oscillation traces a loop around an unstable equilibrium, the point where nullclines intersect. This oscillatory behavior is crucial for modeling biological rhythms, including mitotic Gelens et al. (2014), calcium Sneyd et al. (2017b) and cardiac oscillators Franzone et al. (2006), underscoring the FHN model’s applicability in simulating fast transitions between active and quiescent states in various biological systems.
@@
Excitability. This behavior is intrinsically linked to the oscillatory phenomena mentioned above and pertains to the system’s response to the perturbation of a fixed point. The magnitude of the perturbation determines the system’s response. In the FHN framework, excitability is for example found for parameters $a=0.1$ and $b=1.5$, as illustrated in Fig. 5A.2. Minor perturbations result in a swift return to the baseline state. Conversely, large perturbations drive the system along a nontrivial trajectory (excursion) in the phase space before resettling to the resting state. This trajectory, associated with a ‘spike’ in temporal dynamics, is invariant to the perturbation’s specifics. Excitability is the most used regime in neuroscience and it can reproduce neuronal spikes in the non-coupled cases Parra-Rivas et al. (2023a); FitzHugh (1961); Izhikevich (2006) or in coupled cases Nagumo et al. (1962).
@@
Figure 5: Dynamics and bifurcations of the FHN model for $\varepsilon=0.01$. A. Representative time series and phase space dynamics showing relaxation oscillations ($b=0.5$), excitability ($b=1.5$ and $a=0.1$), and bistability ($b=2$). B. Analytically derived Hopf and saddle node bifurcations, along with homoclinic bifurcations determined numerically, delineating various dynamical regions in the ($a,b$) parameter space. An expanded view of a specific area, marked by a black rectangle, is shown on the right. Black dots correspond to parameters used in A, and dashed lines mark sections analyzed in C. C Bifurcation diagrams of three representative regions. The diagram on the left (C.1) hints at an imminent homoclinic bifurcation due to the close proximity of the limit cycle to the saddle-node. The right diagram (C.3), on the other hand, reveals a shift from stability to instability, lacking an associated limit cycle, suggesting a global bifurcation has taken place. The extrema of limit cycle oscillations are shown by the solid gray line.
Bistability. A nonlinear system exhibits bistability when two stable fixed points coexist for the same set of parameters. This perhaps less frequently examined behavior within the FHN model is depicted in Fig. 5A.3 for parameters $a=0$ and $b=2$. The phase space diagram illustrates how different initial conditions converge to separate stable fixed points, located where the nullclines intersect. Bistability contributes to the emergence of phenomena such as relaxation oscillations, excitability, and hysteresis Boissonade and De Kepper (1980); Rosanov (2002); Hock et al. (1993), vital for the model’s relevance and applicability.
@@
The complexity of the dynamics within the FHN model requires a more detailed analysis, as depicted in Fig. 5B. By examining cross-sections of this figure at constant $b$ values, we can generate 1D bifurcation diagrams, such as those shown in Fig. 5C, plotting the variable $u$ against $a$ for selected $b$ values.
@@
At $b=0.5$ (Fig. 5C.2), a single fixed point that varies monotonically with $a$ is present. This point undergoes Hopf bifurcations at H+ and H-, as indicated in Fig. 5B, leading to (in)stability of the fixed point as shown in Fig. 5C.2, where solid and dashed lines denote stable and unstable solutions, respectively. Unstable fixed points evolve into relaxation oscillations as illustrated in Fig. 5A.1, with limit cycles forming and disappearing at H+ and H-. Through path continuation techniques, we can trace these cycles and assess their stability Doedel et al. (1991). In this parameter regime, oscillations are stable and exhibit amplitude variations as seen in Fig. 5C.2. Close to H+ and H- the oscillations undergo a canard explosion characterized by the abrupt increase of the cycle amplitude for small changes of $a$ Bold et al. (2003).
@@
As $b$ increases, a cusp bifurcation at $(a,b)=(0,2)$ gives rise to three new equilibria: $u_{\text{FP}}^{b}$, $u_{\text{FP}}^{m}$, and $u_{\text{FP}}^{t}$. Fig. 5C.3 for $b=2$ illustrates this, where H+ and H-, previously associated with opposite $a$ values, now occur close to SNr and SNl, respectively. This diagram corresponds to the most right vertical dashed line shown in the close-up view of Fig. 5B. Here the top and bottom equilibrium branches ($u_{\text{FP}}^{t,b}$) are stable nodes and therefore yield a bistable regime. The former stable limit cycle has vanished due to a Saddle-Node bifurcation of Limit Cycles (SNLC) occurring close to H+ and H- (not shown here). At smaller $\varepsilon$ values, the SNLC occurs near H+ and H-, these bifurcations and the Bautin points from which they arise are thoroughly explored in Rocsoreanu’s work Rocsoreanu et al. (2012). Additionally, an unstable limit cycle, emerging from the inverted H+ and H-, quickly vanishes due to another homoclinic bifurcation occurring close to the Hopf bifurcation. For a deeper analysis of these phenomena at larger $\varepsilon$ values, Rocsoreanu et al.’s book offers comprehensive insights Rocsoreanu et al. (2012). Adjacent to this, type II excitability emerges, akin to that shown in Fig. 5A.2, characterized by its proximity to Hopf/SNLC bifurcations, facilitated by the time scale separation and the influence of the former limit cycle in the phase space. As mentioned in the introduction, this type of excitability is characterized by the absence of a clearly defined threshold.
@@
For $b<0$, the system’s bifurcation scenario changes significantly. An illustration for $b=-0.5$ is provided in Fig. 5C.1, where three fixed points persist, albeit with an opposite slope compared to the scenario in Fig. 5C.3. Unlike the bistable regime observed for $b>1$, the top and bottom fixed points ($u_{\text{FP}}^{t,b}$) become unstable saddle points. Additionally, Hopf bifurcations that initially occurred at $u_{\text{FP}}^{t,b}$ now emerge in the middle branch $u_{\text{FP}}^{m}$, rendering it stable between SNl and H-, as well as between H+ and SNr. The relaxation oscillations stemming from H± also experience a canard explosion. Further decreasing $b$ leads to a homoclinic bifurcation (Hom. in Fig. 5B), where the oscillation period diverges, potentially exhibiting type-I excitability Izhikevich (2007); Yelo-Sarrión et al. (2022). This aspect remains largely unexplored within the FHN model.
@@
    Monostable regime with a single fixed point. Defined by a single stable node for $b>0$, this area splits into sectors I.1 and I.2 based on $a$’s positive and negative values, respectively. Both sectors exhibit type I excitability near the Hopf bifurcation, particularly in the presence of canard explosions Izhikevich (2007).
@@
    Bistable regime. Situated between the saddle-node bifurcations SNl,r, this zone’s boundaries are demarcated by H+ for $a>0$ and H- for $a<0$, featuring three equilibria: two stable nodes ($u_{\text{FP}}^{b,t}$) and one saddle point ($u_{\text{FP}}^{m}$).
@@
    Oscillatory regime. This area is known for relaxation oscillations and is further divided into three subregions. In III.1 (for $b<0$), it is confined by Hopf bifurcations H± and the homoclinic bifurcation Hom.; in III.2 (for $b>0$), oscillations are bounded by H+,- and cease after crossing the SNLC, which occurs near the Hopf bifurcation before the intersection of H+ and H-. Canard explosions are observable near the Hopf bifurcations in each subregion.
@@
    Monostable regime within a tri-valued region. For $b<0$, this category includes two monostable regimes where a single stable node ($u_{\text{FP}}^{m}$) coexists with two saddle points ($u_{\text{FP}}^{b,t}$). These settings are located between SNl and H+ (IV.1) and between H- and SNr (IV.2). For $b>0$, a shift from bistable to monostable occurs due to H+ (for $a>0$) and H- (for $a<0$), leading to thin monostable strips flanked by H+ and SNr (for $a<0$) or by H- and SNl (for $a>0$).
@@
    Divergent dynamics regime. Characterized by divergent dynamics due to the lack of an attractor, this regime is prominent for $b<0$ and includes three main divergent sectors: V.1, below SNl, with only the saddle point $u_{\text{FP}}^{b}$; V.2, above SNr, with just the saddle point $u_{\text{FP}}^{t}$; and V.3, within the bounds of Hom., H+, and H-, where the unstable node $u_{\text{FP}}^{m}$ coexists with saddles $u_{\text{FP}}^{b,t}$. Proximity to the homoclinic bifurcation may induce type I excitability Izhikevich (2007).
Figure 6: Time scale separation influences the shape and amplitude of the limit cycle. A. Effect of $\varepsilon$ on the dynamical behavior in phase space and time series for parameters $(a,b)=(0,0.5)$. Setting $\varepsilon=0.5$ leads to less time scale separation and more sinusoidal oscillations. B. Relaxation oscillations disappear for larger values of $a$ and $\varepsilon$. The oscillation amplitude $A(a,\varepsilon)$ becomes dependent on the $a$ parameter, evidenced by the density plot in the left panel and the dispersion of scatter points in the right panel. Furthermore, an inverse relationship between amplitude and $\varepsilon$ is noted, with a decreasing trend in amplitude as $\varepsilon$ increases, indicated by the median and the 50% interquartile range values.
Previously introduced yet not fully explained phenomena, including canard explosions, homoclinic bifurcations (Hom.), and saddle-node bifurcations of limit cycles (SNLC), warrant further exploration to better understand their roles in the dynamics observed.
A canard explosion refers to a quick increase in the limit cycle’s amplitude as the control parameter increases, generating a cycle that occupies a substantial portion of the phase space, instead of a small amplitude oscillation Wechselberger (2007); Benoit (1990). This phenomenon, illustrated in Figs. 5A.1, 5C.1, and 5C.2, is also a characteristic of the Hodgkin-Huxley model Hasan et al. (2018); Rubin and Wechselberger (2007); Moehlis (2006); Ginoux and Llibre (2015) and results from the separation of time scales, a property preserved in the FHN model Rotstein et al. (2012); Parra-Rivas et al. (2023a); Ginoux and Llibre (2015); Rocsoreanu et al. (2012). The resilience of canard cycle amplitudes to parameter perturbations underscores their significance, with the presence of canards signaling type II excitability – a persistence of trajectory behavior after a limit cycle disappears.